Custom Query (419 matches)

Filters
 
Or
 
  
 
Columns

Show under each result:


Results (1 - 3 of 419)

1 2 3 4 5 6 7 8 9 10 11
Ticket Resolution Summary Owner Reporter
#650 fixed Fenix Calibration, end of season, october 2018 asm
Description

This ticket records calibration of the Fenix sensor for the final season. As a memento, the Fenix sensor was repaired by Specim in February 2018 and the detector array was replaced with a new one so timeline comparisons should not be made with calibrations made before 2018

Data location: ~arsf/arsf_data/2018/misc/calibration_october/original_data

#635 fixed GB18/56, flight day 184b/2018, Harwell dac
Description

Data location: /users/rsg/arsf/arsf_data/2018/flight_data/uk/GB18_56-2018_184b_Harwell

Data arrived from NERC-ARF hard drive on 13/07/2018

Scientific objective: Investigate well-defined local methane emission sources in the UK to apply and evaluate the hyperspectral processing method currently developed at the University of Leicester.

PI: Hartmut Boesch BAS Project Code: D056

Issues identified from unpacking checks:

  • Owl capture data missing dark frames

Sensors:

  • Fenix (requested, flown)
  • OWL (requested, flown)
  • Phase One (not requested, flown)
#632 fixed CEH18/16, flight day 177b/2018, Monks Wood dac
Description

Data location: ~arsf/arsf_data/2018/flight_data/uk/CEH18_16-2018_177b_Monks_Wood

Data arrived from NERC-ARF hard drive on 13/07/2018

Scientific objective: Flight over Monks Wood

Priority: Unknown

PI: France Gerard BAS Project Code: C016

The following issues were identified during unpacking

  • Owl data missing GPS start time.

Sensors:

  • Fenix (flown, required)
  • OWL (not flown, required)
  • Phase One (flown, required)
1 2 3 4 5 6 7 8 9 10 11
Note: See TracQuery for help on using queries.